In the old days, folks who went camping used “lantern batteries.”1 Some people still
do, although there is less need now with efficient LED lights. A lantern battery has an
emf of 6.00 V and the bulb in the lantern has a resistance of 20.0 Ω. If the current in
this bulb is 290 mA, what is the internal resistance of the battery? Assume that the
lantern consists solely of the battery and the bulb.
1
Expert's answer
2017-03-03T15:05:05-0500
According to the Ohm’s Law:
I=E/(R+r)
r=E/I-R=(6 V)/(0.29 A)-20 Ω=0.69 Ω